OpenStudy (anonymous):

write the equation that describes the line in slope intercept form. slope=4 point (3,-2)

Vocaloid (vocaloid):

well, we can start with the equation y-y1 = m(x-x1) where m = slope and (x1,y1) is the point (3,-2) plug in your numbers and tell me what you get

OpenStudy (anonymous):

-2-3=4x

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

uh, not quite (x1,y1) = (3,-2) means x1 = 3 and y1 = -2, and we leave x and y alone, giving us: y+2 = 4(x-3) with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

now we convert this to slope intercept form (y = mx + b) y+2 = 4(x-3) start by distributing the 4 y + 2 = 8x - 12 know how to finish from here?

OpenStudy (anonymous):

no @Vocaloid

Vocaloid (vocaloid):

isolate y subtract 2 from each side

OpenStudy (anonymous):

y=8x-10

OpenStudy (anonymous):

is the answer -5/4

Vocaloid (vocaloid):

nope -12 - 2 = ?

OpenStudy (anonymous):

oo my bad -14

Vocaloid (vocaloid):

so your final answer is y = 8x - 14 you are not supposed to solve for y in this scenario

OpenStudy (anonymous):

ooo ok thank you so much :)

OpenStudy (anonymous):

@Vocaloid